oregonduck (oregonduck):

HELP WITH FLVS ALGEBRA 1

oregonduck (oregonduck):

The function H(t) = -16t2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points) Part B: What is the maximum height that the projectile will reach? Show your work. (2 points) Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? (2 points)

oregonduck (oregonduck):

@aouefbaowbu

OpenStudy (mrnood):

part a The projectile hits the ground when H(t) =0 (i.e. its height = 0ft) so put H(t) = 0 in the equation. You will see it is a standard quadratic - so use whatever method you want to solve it . (there will be 2 answers - but one will be evidently impossible.

oregonduck (oregonduck):

ok

oregonduck (oregonduck):

part b?

OpenStudy (mrnood):

do you know differential calculus?

oregonduck (oregonduck):

no this is algebra 1

oregonduck (oregonduck):

A. h(t)=-16t^2+vt+s h(t)=-16t^2+(80)t+(96) h(t)=-16t^2+80t+96

OpenStudy (mrnood):

that is th ecorrect equation . set it =o to solve for time when hits ground

oregonduck (oregonduck):

mine is the correct one?

OpenStudy (mrnood):

yes - but you need to solve it for H(t)=0

oregonduck (oregonduck):

oh ok

OpenStudy (mrnood):

so without calculus for part b:  do you know how to find the vertex for a quadratic equation?